# Bifurcation sequences in the secular 3D planetary 3-Body problem: a geometric approach

### v2

@@ 3.4 Analytical formulas
In the present section we implement our geometric method to analyze the sequence of bifurcations produced in the case in which the planetary system has a small AMD value, i.e. the planetary orbits (2D or 3D) never become very inclined and very eccentric. This allows to adopt a low-order truncation model for the secular Hamiltonian, written generically in Hopf variables as
$\mathcal{K}_{I}(\sigma_{1},\sigma_{3};\sigma_{0})=A\sigma_{1}^{2}+C\sigma_{3}^{2}+B\sigma_{1}\sigma_{3}+(D_{1}\sigma_{0}+\tilde{\Delta}_{1})\sigma_{1}+(D_{3}\sigma_{0}+\tilde{\Delta}_{3})\sigma_{3}\leavevmode\nobreak\ .$ (34)
Without loss of generality, through a rotation we can set $B=0\,$, hence getting the following generic Hamiltonian
$\mathcal{K}_{I}(\sigma_{1},\sigma_{3};\sigma_{0})=\mathcal{A}\sigma_{1}^{2}+\mathcal{C}\sigma_{3}^{2}+(\mathcal{D}_{1}\sigma_{0}+\Delta_{1})\sigma_{1}+(\mathcal{D}_{3}\sigma_{0}+\Delta_{3})\sigma_{3}$ (35)
@@
$\alpha^{2}+\beta^{2}=1\qquad\mathrm{and}\qquad B\alpha^{2}+2A\alpha\beta-2C\alpha\beta-B\beta^{2}=0\leavevmode\nobreak\ \leavevmode\nobreak\ .$
Resulting from a rotation, the coordinates $(\tilde{\sigma}_{0},\tilde{\sigma}_{1},\tilde{\sigma}_{2},\tilde{\sigma}_{3})$ still satisfy the condition of sphere (27) as well as the Poisson algebra $\{\tilde{\sigma}_{i},\tilde{\sigma}_{j}\}=-2\epsilon_{ijk}\tilde{\sigma}_{k}$, $i,j,k=1,2,3\,$. Moreover $\sigma_{0}=\tilde{\sigma}_{0}\,$ is also a Casimir invariant of the algebra in the new variables. Hence, computing the values of $\tilde{\sigma}_{0}$ for which the CP occurs (i.e, solving Eqs. (40) and (41)) is equivalent to finding the radius $\sigma_{0}$ for which the sphere and the surface $\mathcal{C}=\mathcal{E}$ are tangent. For the sake of simplicity, in the following examples we perform the rotation (36) whenever needed and then simply rename the variables $(\tilde{\sigma}_{0},\tilde{\sigma}_{1},\tilde{\sigma}_{2},\tilde{\sigma}_{3})$ in Eq. (35) as $({\sigma}_{0},{\sigma}_{1},{\sigma}_{2},{\sigma}_{3})$.
@@
To find the non-degenerate critical points of the type CP1 of the Hamiltonian (35) we solve Eq.(31), with the assumption that $(\mathcal{D}_{1}\sigma_{0}+\Delta_{1})$ and $(\mathcal{D}_{3}\sigma_{0}+\Delta_{3})$ are not both vanishing. This gives
$\sigma_{1}^{(CP1)}(\mu,\sigma_{0})$ $=$ $-\frac{\mathcal{D}_{1}\sigma_{0}+\Delta_{1}}{2(\mathcal{A}-\mu)}\,,$
$\sigma_{2}^{(CP1)}(\mu,\sigma_{0})$ $=$ $0\,,$ (37)
$\sigma_{3}^{(CP1)}(\mu,\sigma_{0})$ $=$ $-\frac{\mathcal{D}_{3}\sigma_{0}+\Delta_{3}}{2(\mathcal{C}-\mu)}\,.$
Implementing the constraint
${\mathcal{S}}(\mu,\sigma_{0})=(\sigma_{1}^{(CP1)}(\mu,\sigma_{0}))^{2}+(\sigma_{3}^{(CP1)}(\mu,\sigma_{0}))^{2}-\sigma_{0}^{2}=0\,,$
we can look for the conditions to have more than two solutions (in $\mu$) as an effect of changing $\sigma_{0}$ [7]. The numerator of ${\cal S}(\mu,\sigma_{0})=0$ corresponds to an equation of 4th degree in $\mu$, namely
@@
$\mathcal{T}_{1}(\sigma_{0})=\left(\mathcal{D}_{1}+\frac{\Delta_{1}}{\sigma_{0}}\right)^{2},\;\;\mathcal{T}_{3}(\sigma_{0})=\left(\mathcal{D}_{3}+\frac{\Delta_{3}}{\sigma_{0}}\right)^{2}\,.$
@@
$\mathcal{Q}(\sigma_{0})$ $=64(\mathcal{A}-\mathcal{C})^{2}\,\mathcal{T}_{1}(\sigma_{0})\,\mathcal{T}_{3}(\sigma_{0})\Big{[}\left(4(\mathcal{A}-\mathcal{C})^{2}-\mathcal{T}_{1}(\sigma_{0})\right)^{3}-3\mathcal{T}_{3}(\sigma_{0})\Big{(}16(\mathcal{A}-\mathcal{C})^{4}$ (39)
$\phantom{=}+28\mathcal{T}_{1}(\sigma_{0})(\mathcal{A}-\mathcal{C})^{2}+\mathcal{T}_{1}(\sigma_{0})^{2}\Big{)}+3\mathcal{T}_{3}(\sigma_{0})^{2}\left(4(\mathcal{A}-\mathcal{C})^{2}-\mathcal{T}_{1}(\sigma_{0})\right)-\mathcal{T}_{3}(\sigma_{0})^{3}\Big{]}$
Notice that $\mathcal{T}_{1}(\sigma_{0})=0$ or $\mathcal{T}_{3}(\sigma_{0})=0$ do not correspond to a change in the sign of the discriminant. At $\mathcal{A}=\mathcal{C}$ the discriminant is identically zero for every value of $\sigma_{0}$; this is the situation at which the ellipses degenerate to circles. Therefore, tangencies with the phase space can occur either at an infinite number of points or at no point. Higher order normal forms would be needed to describe the bifurcations. We do not consider this case here. By solving $\mathcal{Q}(\sigma_{0})=0$ with respect to $\mathcal{T}_{3}(\sigma_{0})$ and considering the only real solution, the bifurcation values of $\sigma_{0}$ can be found by imposing
$f_{1}(\sigma_{0})=0$
@@
Instead, the bifurcation values of $\sigma_{0}$ for fixed points of the type CP2 can be found by solving
$f_{2}(\sigma_{0})=-4+\frac{\mathcal{T}_{1}(\sigma_{0})}{\mathcal{A}^{2}}+\frac{\mathcal{T}_{3}(\sigma_{0})}{\mathcal{C}^{2}}=0\,.$
The solution in this case can be explicitly found:
$\begin{split}&\sigma_{0}^{(CP2,1)}=-\frac{\mathcal{C}^{2}\mathcal{D}_{1}\Delta_{1}+\mathcal{A}^{2}\mathcal{D}_{3}\Delta_{3}+\mathcal{A}\mathcal{C}\sqrt{4\mathcal{C}^{2}\Delta_{1}^{2}-(\mathcal{D}_{3}\Delta_{1}+2\mathcal{A}\Delta_{3}-\mathcal{D}_{1}\Delta_{3})(\mathcal{D}_{3}\Delta_{1}-(2\mathcal{A}+\mathcal{D}_{1})\Delta_{3})}}{\mathcal{C}^{2}\mathcal{D}_{1}^{2}+\mathcal{A}^{2}(-4\mathcal{C}^{2}+\mathcal{D}_{3}^{2})}\,,\\ &\sigma_{0}^{(CP2,2)}=-\frac{\mathcal{C}^{2}\mathcal{D}_{1}\Delta_{1}+\mathcal{A}^{2}\mathcal{D}_{3}\Delta_{3}-\mathcal{A}\mathcal{C}\sqrt{4\mathcal{C}^{2}\Delta_{1}^{2}-(\mathcal{D}_{3}\Delta_{1}+2\mathcal{A}\Delta_{3}-\mathcal{D}_{1}\Delta_{3})(\mathcal{D}_{3}\Delta_{1}-(2\mathcal{A}+\mathcal{D}_{1})\Delta_{3})}}{\mathcal{C}^{2}\mathcal{D}_{1}^{2}+\mathcal{A}^{2}(-4\mathcal{C}^{2}+\mathcal{D}_{3}^{2})}\,.\end{split}$ (41)
Geometrically, the appearance of the CP2 is due to the centers of the ellipses or hyperbolas entering or leaving the circle $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}:\sigma_{1}^{2}+\sigma_{3}^{2}=\sigma_{0}^{2}$.

### v3

@@ 3.4 Analytical formulas
In the present section we implement our geometric method to analyze the sequence of bifurcations produced in the case in which the planetary system has a small AMD value, i.e. the planetary orbits (2D or 3D) never become very inclined and very eccentric. In this case, we can use a low-order truncation of the Hamiltonian model, leading to simplified expressions which can be treated by explicit analytical formulas. Note that the issue of the proper truncation of the Hamiltonian has been studied as regards i) the maximum multipole order ([6]), or ii) the truncation order in the orbital eccentricities and inclinations ([8]). Furthermore the accuracy of the integrable approximation with respect to the full secular Hamiltonian has been studied in subsection 3.2 of [17].
More specifically, we consider below the low-order truncation model of the secular Hamiltonian written generically in Hopf variables as
$\mathcal{K}_{I}(\sigma_{1},\sigma_{3};\sigma_{0})=A\sigma_{1}^{2}+C\sigma_{3}^{2}+B\sigma_{1}\sigma_{3}+(D_{1}\sigma_{0}+\tilde{\Delta}_{1})\sigma_{1}+(D_{3}\sigma_{0}+\tilde{\Delta}_{3})\sigma_{3}\penalty 10000\ .$ (38)
Without loss of generality, through a rotation we can set B=0B=0\,, hence getting the following generic Hamiltonian
$\mathcal{K}_{I}({\tilde{\sigma}_{1},\tilde{\sigma}_{3};\tilde{\sigma}_{0}})=\mathcal{A}{\tilde{\sigma}_{1}}^{2}+\mathcal{C}{\tilde{\sigma}_{3}}^{2}+(\mathcal{D}_{1}{\tilde{\sigma}_{0}}+\Delta_{1}){\tilde{\sigma}_{1}}+(\mathcal{D}_{3}{\tilde{\sigma}_{0}}+\Delta_{3}){\tilde{\sigma}_{3}}$ (39)
@@
$\alpha^{2}+\beta^{2}=1\qquad\mathrm{and}\qquad B\alpha^{2}+2A\alpha\beta-2C\alpha\beta-B\beta^{2}=0\penalty 10000\ \penalty 10000\ .$
Resulting from a rotation, the coordinates $(\tilde{\sigma}_{0},\tilde{\sigma}_{1},\tilde{\sigma}_{2},\tilde{\sigma}_{3})$ still satisfy the condition given by Eq. (31), i.e., $(\tilde{\sigma}_{1},\tilde{\sigma}_{2},\tilde{\sigma}_{3})$ are on a sphere of radius $\tilde{\sigma}_{0}$, as well as the Poisson algebra $\{\tilde{\sigma}_{i},\tilde{\sigma}_{j}\}=-2\epsilon_{ijk}\tilde{\sigma}_{k}$, i,j,k=1,2,3i,j,k=1,2,3\,. Moreover $\sigma_{0}=\tilde{\sigma}_{0}\,$ is also a Casimir invariant of the algebra in the new variables. Hence, computing the values of $\tilde{\sigma}_{0}$ for which the CP occurs (i.e, solving Eqs. (47) and (50)) is equivalent to finding the radius $\sigma_{0}$ for which the sphere and the surface $\mathcal{C}_{\sigma_{0},\mathcal{E}}$} are tangent. For the sake of simplicity, in the following examples we perform the rotation (40) whenever needed and then simply rename the variables $(\tilde{\sigma}_{0},\tilde{\sigma}_{1},\tilde{\sigma}_{2},\tilde{\sigma}_{3})$ in Eq. (39) as $({\sigma}_{0},{\sigma}_{1},{\sigma}_{2},{\sigma}_{3})$.
@@
To find the non-degenerate critical points of the type CPI of the Hamiltonian (39) we solve Eq. (35), with the assumption that $(\mathcal{D}_{1}\sigma_{0}+\Delta_{1})$ and $(\mathcal{D}_{3}\sigma_{0}+\Delta_{3})$ are not vanishing. This gives
$\sigma_{1}^{({CPI})}(\mu,\sigma_{0})$ $=$ $-\frac{\mathcal{D}_{1}\sigma_{0}+\Delta_{1}}{2(\mathcal{A}-\mu)}\,$,
$\sigma_{2}^{({CPI})}(\mu,\sigma_{0})$ $=$ $0\,$, (41)
$\sigma_{3}^{({CPI})}(\mu,\sigma_{0})$ $=$ $-\frac{\mathcal{D}_{3}\sigma_{0}+\Delta_{3}}{2(\mathcal{C}-\mu)}\,.$
If $(\mathcal{D}_{1}\sigma_{0}+\Delta_{1})=0$, Hamiltionian (39) becomes symmetric with respect to the reflection $\sigma_{1}\rightarrow-\sigma_{1}$, corresponding to the $2:2$ resonance (see [16]). Then equation (35) is solved by
$\mu=\mathcal{A}\;\;,\,\,\sigma_{3}^{({CPI})}(\mu,\sigma_{0})=-\frac{\mathcal{D}_{3}\sigma_{0}+\Delta_{3}}{2(\mathcal{C}-\mu)}.$ (42)
Similarly, if $(\mathcal{D}_{3}\sigma_{0}+\Delta_{3})=0$ we find the solution to be
$\mu=\mathcal{C}\;\;,\,\,\sigma_{1}^{({CPI})}(\mu,\sigma_{0})=-\frac{\mathcal{D}_{1}\sigma_{0}+\Delta_{1}}{2(\mathcal{A}-\mu)}.$ (43)
Exchanging $\sigma_{3}\rightarrow\sigma_{1}$, the case $(\mathcal{D}_{3}\sigma_{0}+\Delta_{3})=0$ corresponds to $(\mathcal{D}_{1}\sigma_{0}+\Delta_{1})=0$. The bifurcation sequences of the Hamiltonian (39) are well understood in this last case (see [16]). In the present work, instead, we focus on the study of the generic case in which $(\mathcal{D}_{j}\sigma_{0}+\Delta_{j})\neq 0$, $j=1,3$.
3.4.1 Computation of the bifurcation value of $\sigma_{0}$
The value of $\mu$ corresponding to the solutions of Eq. (35) is given by the constraint
${\mathcal{S}}(\mu,\sigma_{0})=(\sigma_{1}^{({CPI})}(\mu,\sigma_{0}))^{2}+(\sigma_{3}^{({CPI})}(\mu,\sigma_{0}))^{2}-\sigma_{0}^{2}=0\,$,
where the expressions for $\sigma_{1}^{({CPI})}$ and $\sigma_{3}^{({CPI})}$ are provided in Eq. (3.4). There exist at least two solutions to the above equations for $\mu$ (see [9]). Thus, we look for the critical values of $\sigma_{0}$ determining a change in the number of real solutions in $\mu$ from $2$ to $4$, or viceversa.
The numerator of ${\cal S}(\mu,\sigma_{0})=0$ corresponds to an equation of 4th degree in $\mu$, namely
@@ 3.4.1 Computation of the bifurcation value of $\sigma_{0}$
$\mathcal{T}_{1}(\sigma_{0})=\left(\mathcal{D}_{1}+\frac{\Delta_{1}}{\sigma_{0}}\right)^{2},\;\;\mathcal{T}_{3}(\sigma_{0})=\left(\mathcal{D}_{3}+\frac{\Delta_{3}}{\sigma_{0}}\right)^{2}\,.$ (45)
@@
[…]
Notice that the values of $\sigma_{0}$ such that $\mathcal{T}_{1}(\sigma_{0})=0$ or $\mathcal{T}_{3}(\sigma_{0})=0$ do not correspond to a change in the sign of the discriminant , since $\mathcal{T}_{1}(\sigma_{0})$ and $\mathcal{T}_{3}(\sigma_{0})$ are quadratic expressions. At $\mathcal{A}=\mathcal{C}$ the discriminant is identically zero for every value of $\sigma_{0}$; this is the situation at which the ellipses degenerate to circles. Therefore, tangencies with the phase space can occur either at an infinite number of points or at no point. Higher order normal forms would be needed to describe the bifurcations. We do not consider this case here, which, anyway, has no particular physical significance and would appear only through a coincidence of the values of some Laplace coefficients, for finely tuned values of the planetary masses and/or semi-major axes.
Dividing Eq. (46) by $\mathcal{T}_{3}(\sigma_{0})$, the discriminant with respect to $\mathcal{T}_{3}(\sigma_{0})$ turns to be always negative. This means that, apart from the solutions $\mathcal{T}_{3}(\sigma_{0})=0$, the equation $\mathcal{Q}(\sigma_{0})=0$ has only one real solution for $\mathcal{T}_{3}(\sigma_{0})$, namely
$\mathcal{T}_{3}(\sigma_{0})=4(\mathcal{A}-\mathcal{C})^{2}-\mathcal{T}_{1}(\sigma_{0})+2^{2/3}\,3\Big((\mathcal{A}-\mathcal{C})^{2}\,\mathcal{T}_{1}(\sigma_{0})^{2}\Big)^{1/3}-2^{1/3}\,6\Big((\mathcal{A}-\mathcal{C})^{4}\,\mathcal{T}_{1}(\sigma_{0})\Big)^{1/3}\,.$
Recalling the definitions of $\mathcal{T}_{1}(\sigma_{0})$ and $\mathcal{T}_{3}(\sigma_{0})$ (Eq. (45)), it is easy to see that the above expression gives an equation for the bifurcation values of $\sigma_{0}$, namely
$f_{1}(\sigma_{0})=0\,$, (47)
@@
$f_{1}(\sigma_{0})=-4(\mathcal{A}-\mathcal{C})^{2}+\mathcal{T}_{1}(\sigma_{0})+\mathcal{T}_{3}(\sigma_{0})-2^{2/3}\,3\Big((\mathcal{A}-\mathcal{C})^{2}\,\mathcal{T}_{1}(\sigma_{0})^{2}\Big)^{1/3}+2^{1/3}\,6\Big((\mathcal{A}-\mathcal{C})^{4}\,\mathcal{T}_{1}(\sigma_{0})\Big)^{1/3}\,.$ (48)
For real not negative roots $\sigma_{0}$ of Eq. (47), the discriminant with respect to $\mu$ of Eq. (44) (i.e. Eq. (46)) vanishes. This implies that a pair of real solutions in $\mu$ of Eq. (44) appears/disappears. Correspondingly, a pair of equilibria appears/disappears for the Hamiltonian (39), i.e., we obtain a bifurcation. In general, we cannot predict the number of non-negative real solutions of Eq. (47) for $\sigma_{0}$, as this depends on the specific values of $\mathcal{A},\mathcal{C},\mathcal{D}_{j},\Delta_{j}$. However, since Eq. (44) is of fourth degree in $\mu$ and we know that at least two real solutions for $\mu$ always exist ([9]), we can conclude that for the values of $\sigma_{0}$ that solve Eq. (47), the number of fixed points of the type CPI of system (39) changes from $2$ to $4$, or vice versa.
Instead, if $\mathcal{A}\mathcal{C}\neq 0$, the bifurcation values of $\sigma_{0}$ for fixed points of the type CPII can be found by solving
$f_{2}(\sigma_{0})=-4+\frac{\mathcal{T}_{1}(\sigma_{0})}{\mathcal{A}^{2}}+\frac{\mathcal{T}_{3}(\sigma_{0})}{\mathcal{C}^{2}}=0\,.$ (49)
In the case $\mathcal{A}=0$ or $\mathcal{C}=0$, there are no isolated equilibria of the second kind (i.e. CPII). This is a degenerate case that we do not consider in this work. If $\mathcal{C}^{2}\mathcal{D}_{1}^{2}+\mathcal{A}^{2}(-4\mathcal{C}^{2}+\mathcal{D}_{3}^{2})\neq 0$, we find two solutions for Eq. (49), given by
@@
provided that
$4\mathcal{C}^{2}\Delta_{1}^{2}-(\mathcal{D}_{3}\Delta_{1}+2\mathcal{A}\Delta_{3}-\mathcal{D}_{1}\Delta_{3})(\mathcal{D}_{3}\Delta_{1}-(2\mathcal{A}+\mathcal{D}_{1})\Delta_{3})\geq 0.$
Instead, we find
$\sigma_{0}^{({CPII})}=-\frac{(\mathcal{C}^{2}\Delta_{1}^{2}+\mathcal{A}^{2}\Delta_{3}^{2})|\mathcal{C}|}{2(\mathcal{A}^{2}\mathcal{D}_{3}\Delta_{3}|\mathcal{C}|\pm\mathcal{C}^{2}\Delta_{1}|\mathcal{A}|\sqrt{4\mathcal{C}^{2}-\mathcal{D}_{3}^{2})}}\,$ (51)
if $\mathcal{D}_{1}=\pm\frac{|\mathcal{A}|}{|\mathcal{C}|}\sqrt{(4\mathcal{C}^{2}-\mathcal{D}_{3}^{2})}$, respectively. In this case, necessarily $\mathcal{A}^{2}\mathcal{D}_{3}\Delta_{3}|\mathcal{C}|\pm\mathcal{C}^{2}\Delta_{1}|\mathcal{A}|\sqrt{4\mathcal{C}^{2}-\mathcal{D}_{3}^{2}}\neq 0$ and $4\mathcal{C}^{2}-\mathcal{D}_{3}^{2}\geq 0$. Otherwise if $\mathcal{A}^{2}\mathcal{D}_{3}\Delta_{3}|\mathcal{C}|\pm\mathcal{C}^{2}\Delta_{1}|\mathcal{A}|\sqrt{4\mathcal{C}^{2}-\mathcal{D}_{3}^{2}}=0$ there is no solution of Eq. (49) for $\sigma_{0}$.
Geometrically, the fixed points of type CPII correspond to solutions of Eq. (35) at which the surface $\mathcal{C}_{\sigma_{0},\mathcal{E}}$ of constant energy of the Hamiltonian (39) degenerates to a straight line parallel to the $\sigma_{2}$-axis or to two sheets intersecting along a line parallel to the $\sigma_{2}$-axis. At the values of $\sigma_{0}^{(CPII)}$ given by (50) or (51), this line is tangent to the sphere $\mathcal{S}_{\sigma_{0}^{(CPII)}}$. In particular, on the plane $\sigma_{2}=0$, the center of the family of ellipses or hyperbolas corresponding to $\mathcal{C}_{\sigma_{0}^{(CPII)},\,\mathcal{E}}$ lies on the circle $\mathcal{S}_{\sigma_{0}^{(CPII)}}^{(\sigma_{2}=0)}$, i.e. $\sigma_{1}^{2}+\sigma_{3}^{2}=\left(\sigma_{0}^{(CPII)}\right)^{2}.$ The appearance/disappearance of the CPII occurs when the center of the ellipses or hyperbolas enters or leaves this circle (see Section 3.3.2).
@@ 4 Application to different integrable models approximating the secular Hamiltonian
In this section, we apply the method of the previous section in order to compute the sequence of bifurcations in three different Hamiltonian models approximating the extrasolar planetary system, i.e., the Hamiltonians $\mathcal{H}_{int}$, $\mathcal{H}_{int}^{(1)}$ and $\tilde{\mathcal{H}}_{int}^{(N_{P}=3,\,N_{bk}=4)}$. As already mentioned, $\mathcal{H}_{int}$ represents the integrable approximation of the full secular Hamiltonian analyzed in Section 2 (see Eq. (6)). Since we use a secular Hamiltonian model truncated at multipolar order $N_{P}=5$ and of book-keeping order (in eccentricity and inclination) $N_{bk}=12$ (see Section 2.2), the integrable Hamiltonian $\mathcal{H}_{int}$, expressed in the Hopf variables contains powers in $\sigma_{i}$, $i=0,\dots,3$ up to 66-th degree. For this reason, in the analyses of Section 4.1, as well as of Section 4.2, we can apply the general discussion of Sections 3.1-3.3, but we cannot apply the analytical formulæ described in Section 3.4. In fact, the exact equality given in Eq. (38), that is the basis for the development of the analytical formulæ, would be valid just for a Hamiltonian model developed up to fourth order in eccentricity. In the case of $\mathcal{H}_{int}$ and $\mathcal{H}_{int}^{(1)}$, Eq. (38) is just an approximation and not an exact equality. The approximation is sufficient to understand the global behavior of the system and the shape of the curves through which the bifurcation sequences are computed, but not to provide quantitative estimates on the bifurcation values. Instead, the analytical formulæ of subsection 3.4 can be used in Section 4.3, where the equality is valid.